# On the minimax optimality of Flow Matching through the connection to kernel density estimation

### v1

@@
(Karlsruhe Insitute of Technology)
Flow Matching has recently gained attention in generative modeling as a simple and flexible alternative to diffusion models, the current state of the art. While existing statistical guarantees adapt tools from the analysis of diffusion models, we take a different perspective by connecting Flow Matching to kernel density estimation. We first verify that the kernel density estimator matches the optimal rate of convergence in Wasserstein distance up to logarithmic factors, improving existing bounds for the Gaussian kernel. Based on this result, we prove that for sufficiently large networks, Flow Matching also achieves the optimal rate up to logarithmic factors, providing a theoretical foundation for the empirical success of this method. Finally, we provide a first justification of Flow Matching’s effectiveness in high-dimensional settings by showing that rates improve when the target distribution lies on a lower-dimensional linear subspace.
@@ 1 Introduction
Flow Matching, a generative model introduced by Lipman et al. [2023], has recently attained significant interest due to its considerably more straightforward construction in comparison to diffusion models (building up on Sohl-Dickstein et al. [2015]), which have been regarded as the state-of-the-art generative method. Instead of learning $\psi$ directly, like GANs (Goodfellow et al. [2014]), Flow Matching relies on learning a vector field $v$ that describes the direction of movement to transfer the probability mass from $\mathbb{U}$ to $\mathbb{P}^{*}$. This leads to the following ODE in time $t\in[0,1]$,
@@
where $p_{0}$ is the density of $\mathbb{U}$. The vector field $v$ should be chosen such that $p_{1}$ is close to the density of $\mathbb{P}^{*}$.
Originally, normalizing flows were used to transform a latent distribution into an approximation of an unknown distribution by an invertible map constructed according to the change of variable theorem. This method was defined by Tabak and Vanden-Eijnden [2010] and Tabak and Turner [2013], but popularized by Rezende and Mohamed [2015] and Dinh et al. [2015]. For an overview see Kobyzev et al. [2020]. Continuous normalizing flows (CNF, Chen et al. [2018]) use a neural ordinary differential equation (ODE) to construct this mapping via a maximum likelihood method. However, the training requires simulations of the ODE, which results in high computational cost, see Grathwohl et al. [2019]. The Flow Matching approach of Lipman et al. [2023] avoids this by using a regression objective to train a CNF, which avoids simulations during training,
@@
where $p_{t}(\cdot|Y)$ is a time dependent conditional density path associated to the conditional vector field $v_{t}(\cdot|Y)$ and both $p_{t}(\cdot|Y)$ and $v_{t}(\cdot|Y)$ are known by construction. This regression objective is in certain cases similar to the score matching objective (Hyvärinen [2005], Vincent [2011]), which is used to train diffusion models based on stochastic differential equations, introduced by Song et al. [2020].
@@
Flow Matching algorithms have been successfully used in many different applications that benefit from efficient sampling, such as text-to-speech (Guo et al. [2024]) and text-to-image (Yang et al. [2024], Esser et al. [2024]) settings, the production of novel molecular or protein structures (Dunn and Koes [2024], Bose et al. [2024]) or the construction of surrogate models in high energy physics (Bieringer et al. [2024]). It has also been adjusted theoretically to different settings. Atanackovic et al. [2024] adapted Flow Matching to the case of interacting particles, Gat et al. [2024] explored the discrete setting. Chen and Lipman [2024] generalized the Euclidean setting to the Riemannian setting, allowing for more general geometries. Kerrigan et al. [2023] extended Flow Matching to function spaces. However, the statistical properties have only recently been studied by Fukumizu et al. [2024] in the Wasserstein $2$ distance. They do not use the exact setting of Lipman et al. [2023], but rather introduce stopping times that depend on the number of samples, enabling the transfer of methods known from the statistical analysis of diffusion models.
@@
In this work, we demonstrate that Flow Matching in the setting of Lipman et al. [2023] is closely related to the classical kernel density estimation (KDE). This connection allows us to analyze Flow Matching from a new perspective. First, we show that the motivation of Flow Matching also holds for its empirical counterparts where we replace the unknown target density $p^{*}$ in (1) with the empirical measure of the observations. For sufficiently large network classes, the resulting generative algorithm coincides exactly with a kernel density estimator, where the kernel is given by the density of the latent distribution. We show convergence rates for the kernel density estimator in Wasserstein-$1$ distance in a setting that allows for the Gaussian kernel which is the most popular choice in Flow Matching applications. Our rates coincide with the minimax optimal rates from Niles-Weed and Berthet [2022] up to logarithmic factors. We then transfer our analysis to Flow Matching using neural network classes of parameterized vector fields.
@@
Unlike Fukumizu et al. [2024], we do not build on the similarities to diffusion models. Separating the error of the kernel density estimator allows us to use empirical risk minimization without the need for Chernoff-type bounds. this avoids the problems pointed out by Yakovlev and Puchkin [2025] in the analysis of diffusion models. Overall, the analysis of Flow Matching is more delicate, since Girsanov’s theorem does not apply and thus the strategy of Chen et al. [2023c] cannot be used. Therefore, our bound depends exponentially on the Lipschitz constant of the vector field. This is one of the reasons for the difficulties in the proof of Fukumizu et al. [2024]. This analysis allows us to show convergence rates in Wasserstein-$1$ distance in the case where finite neural networks are used for the vector field. While these rates are minimax optimal up to a logarithmic constant, they suffer from the curse of dimensionality. In case the support of $\mathbb{P}^{*}$ is concentrated on a linear subspace, we improve our results in a way that the convergence rate depends only on the intrinsic dimension. This provides a first justification for the excellent performance of Flow Matching for high-dimensional data sets.
@@
Lipman et al. [2023] use a fixed latent distribution. Similar approaches for flows between two possibly unknown distributions $\mathbb{P}$ and $\mathbb{Q}$ are studied by Tong et al. [2024], Liu et al. [2022] and Albergo and Vanden-Eijnden [2022]. Tong et al. [2024] also generalize the mentioned methods. Gao et al. [2024] prove a suboptimal rate of convergence for certain forms of Flow Matching not exactly conforming to the construction on Lipman et al. [2023] in the Wasserstein $2$ distance. Benton et al. [2024] analyzed Flow Matching excluding the approximation error by imposing assumptions on the covariance that lead to global Lipschitz bounds on the vector field. They also focus on different constructions than Lipman et al. [2023]. Gong et al. [2025] study the properties of ReLU networks to approximate a vector field corresponding to higher order trajectories.
As Lipman et al. [2023] point out, Flow Matching is closely related to diffusion models. For an overview of generative diffusion models, see Cao et al. [2024]. Indeed, even in cases that are not constructed to be consistent with diffusion models, the approximation of a score function has similar properties to the approximation of the Flow Matching vector field. Fukumizu et al. [2024] build up on this connection. The statistical properties of score matching are an area of ongoing research, see for example Chen et al. [2023a], Chen et al. [2023b], Chen et al. [2023c], Oko et al. [2023], Tang and Yang [2024], Azangulov et al. [2024], Zhang et al. [2024] , Yakovlev and Puchkin [2025]. Marzouk et al. [2023] study the statistical properties of CNFs trained by likelihood maximization.
The kernel density or Parzen-Rosenblatt estimator (Rosenblatt [1956], Parzen [1962]) is the classical method for estimating a smooth density. For an overview in the univariate case, see Tsybakov [2009], Chapter 1.2 or Devroye and Lugosi [2001] and in the multivariate case see Scott [1992]. Typically, the kernel density estimator is analyzed using the mean squared error or the $L_{1}$ distance. Despite its longstanding use, kernel density estimation is still the subject of ongoing research, for example in density estimation on unknown manifolds (Berry and Sauer [2017], Berenfeld and Hoffmann [2021], Divol [2022], Wu and Wu [2022]). In particular, Divol [2022] uses the Wasserstein metric to evaluate the performance of a kernel density estimator, however he needs to assume properties of the kernel that are not satisfied by standard choices such as the Gaussian kernel.
@@
In Section 2 we introduce the problem setting mathematically and recall all the definitions needed to define the conditional Flow Matching objective from Lipman et al. [2023]. We introduce the empirical counterparts, validate the empirical conditional Flow Matching objective, and portray the connection to the kernel density estimator. In Section 3 we start by separating the error corresponding to the kernel density estimator and the vector field approximation. In Section 3 we first separate the error corresponding to the kernel density estimator and the vector field approximation. We then proceed to bound the former error in Section 3.1. Incorporating theory of network approximation, we bound the latter error and obtain a rate of convergence for the entire Flow Matching procedure in Section 3.2. Subsequently, we show in Section 3.3 that the rate can be improved if the unknown distribution is supported on a linear subspace. All proofs are deferred to Section 4.
@@ 1.1 Some notations
We write $X\sim p$ as a synonym for $X\sim\mathbb{P}$ in case $\mathbb{P}$ has a density $p$ with respect to the Lebesgue measure and $\mathbb{P}$ is the distribution of the random variable $X$. By $\lesssim$ we denote $a\leq c\cdot b,$ where $a,b\in\mathbb{R}$ and $c$ is a constant independent of $n$. For $d\in\mathbb{N}$, we use the Euclidean norm ||$\cdot$|| ⋅ | on $\mathbb{R}^{d}$. The supremum norm and the $L_{1}$-norm of a function $f\colon\Omega\rightarrow\mathbb{R}^{d^{\prime\prime}}$ with $d^{\prime},d^{\prime\prime}\in\mathbb{R}$ and $\Omega\subset\mathbb{R}^{d^{\prime}}$ is defined as
@@
For $d^{\prime\prime}=1$ and $\alpha\in(0,1],$ the Besov space $B_{1,\infty}^{\alpha}$ is defined as
$B_{1,\infty}^{\alpha}(\Omega)\coloneqq\big{\{}f\in L_{1}(\Omega)\colon|f|_{B_{1,\infty}^{\alpha}(\Omega)}\coloneqq\sup_{t>0}t^{-\alpha}\omega_{1}(f,t)_{1}<\infty\big{\}}$
@@
The difference is set to zero if $x+h\notin\Omega.$ On $B_{1,\infty}^{\alpha}$ we define the norm $\|f\|_{B_{1,\infty}^{\alpha}(\Omega)}=\|f\|_{L_{1}(\Omega)}+|f|_{B_{1,\infty}^{\alpha}(\Omega)}$.
@@ 2 Empirical Flow Matching
The ODE (2) corresponds to the Lagrangian description (in terms of particle trajectories) of the conservation of mass formula [Villani, 2008, p.14]. The change of variables formula links it to the Eulerian description: A necessary and sufficient condition for $v_{t}$ to generate $p_{t}$ is
@@
by Lipman et al. [2023]. They have consider a probability path of the form
@@
[…]
In this setting Lipman et al. [2023] show that the Flow Matching objective (4) is equivalent to the conditional Flow Matching objective
@@
With this modification we recover the sufficient condition for $v^{n}_{t}$ to generate $p^{n}_{t}$ analogously to [Lipman et al., 2023, Theorem 1].
@@
The motivation for the conditional Flow Matching objective (7) is the equivalence to the unconditioned Flow Matching objective with respect to the optimizing arguments [Lipman et al., 2023, Theorem 2]. Using the empirical counterparts, this still holds true.
@@
Using $p^{n}_{t}$ and $v^{n}_{t}$ from (8), and a class of parameterized vector fields $\mathcal{M},$ which is constructed such that the minimal arguments exists, we have that
$\underset{\hat{v}\in\mathcal{M}}{\mathrm{argmin}}\int_{0}^{1}\mathbb{E}_{X_{t}\sim p^{n}_{t}}\big{[}|\hat{v}_{t}(X_{t})-v^{n}_{t}(X_{t})|^{2}\big{]}\;\mathrm{d}t=\underset{\hat{v}\in\mathcal{M}}{\mathrm{argmin}}\int_{0}^{1}\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\tilde{X}_{t}\sim p_{t}(\cdot|X_{i}^{*})}\big{[}|\hat{v}_{t}(\tilde{X}_{t})-v_{t}(\tilde{X}_{t}|X_{i}^{*})|^{2}\big{]}\;\mathrm{d}t.$
@@
is justified theoretically. Compared to Gao et al. [2024], we use $v_{t}(\cdot|X_{i}^{*})$ directly as proposed by Lipman et al. [2023] and stick to the entire time interval. Note that for simplicity we do not sample $t.$
From minimizing (9) in $\tilde{v}\in\mathcal{M},$ we obtain an optimal argument $\hat{v}.$ Solving the ODE (2) using $\hat{v}$, we obtain a flow $\hat{\psi}_{t}$, i.e. $\hat{\psi}$ is given by
$\frac{d}{dt}\hat{\psi}_{t}(x)=\hat{v}_{t}(\hat{\psi}_{t}(x)),\quad\hat{\psi}_{0}(x)=x,\qquad\text{for}\quad\hat{v}\in\operatorname*{argmin}_{\tilde{v}\in\mathcal{M}}\tilde{\Psi}(\tilde{v}).$ (10)
@@
we recover the same result like in [Lipman et al., 2023, Theorem 3]:
@@
Set $\sigma_{\min}>0.$ To flow from $p^{n}_{0}(x)=\frac{1}{n}\sum_{i=1}^{n}K(x)=K(x)$ to $p^{n}_{1}(x)=\frac{1}{n\sigma^{d}_{\mathrm{min}}}\sum_{i=1}^{n}K(\frac{x-X_{i}^{*}}{\sigma_{\mathrm{min}}}),$ we can choose any differentiable functions $\sigma_{t}$ and $\mu_{t}$ such that
@@
$p^{n}_{1}(x)=\frac{1}{n\sigma^{d}_{\mathrm{min}}}\sum_{i=1}^{n}K\Big{(}\frac{x-X_{i}^{*}}{\sigma_{\mathrm{min}}}\Big{)},$ (13)
where the kernel is given by the latent distribution $\mathbb{U}$. Choosing $\mathbb{U}=\mathcal{N}_{d}(0,1)$, i.e., we consider the $d$-dimensional Gaussian kernel $K(x)=(2\pi)^{-d/2}\exp(-|x|^{2}/2)$, yields the proposed flow from Lipman et al. [2023, Section 4]. Moreover, considering general kernels is in line with methods by Tong et al. [2024], Liu et al. [2022] and Albergo and Vanden-Eijnden [2022], which transform an unknown distribution to another.

### v2

@@
(Ruhr University Bochum and Karlsruhe Institute of Technology)
Flow Matching has recently gained attention in generative modeling as a simple and flexible alternative to diffusion models. While existing statistical guarantees adapt tools from the analysis of diffusion models, we take a different perspective by connecting Flow Matching to kernel density estimation. We first verify that the kernel density estimator matches the optimal rate of convergence in Wasserstein distance up to logarithmic factors, improving existing bounds for the Gaussian kernel. Based on this result, we prove that for sufficiently large networks, Flow Matching achieves the optimal rate up to logarithmic factors. If the target distribution lies on a lower-dimensional manifold, we show that the kernel density estimator profits from the smaller intrinsic dimension on a small tube around the manifold. The faster rate also applies to Flow Matching, providing a theoretical foundation for its empirical success in high-dimensional settings.
@@ 1 Introduction
Flow Matching, a generative model introduced by Lipman et al. (2023), has recently attained significant interest due to its considerably more straightforward construction in comparison to diffusion models (building up on Sohl-Dickstein et al. (2015)), which have been regarded as the state-of-the-art generative method. Instead of learning $\psi$ directly, like GANs (Goodfellow et al., 2014), Flow Matching relies on learning a vector field $v$ that describes the direction of movement to transfer the probability mass from $\mathbb{U}$ to $\mathbb{P}^{*}$. This leads to the following ODE in time $t\in[0,1]$,
@@
where $p_{0}$ is the density of $\mathbb{U}$. The vector field $v$ should be chosen such that $p_{1}$ is close to the density $p^{*}$ of $\mathbb{P}^{*}$.
Originally, normalizing flows were used to transform a latent distribution into an approximation of an unknown distribution by an invertible map constructed according to the change of variable theorem. This method was defined by Tabak and Vanden-Eijnden (2010) and Tabak and Turner (2013), but popularized by Rezende and Mohamed (2015) and Dinh et al. (2015). For an overview see Kobyzev et al. (2020). Continuous normalizing flows (CNF), introduced by Chen et al. (2018), use a neural ordinary differential equation (ODE) to construct this mapping via a maximum likelihood method. However, the training requires simulations of the ODE, which results in high computational cost, see Grathwohl et al. (2019). The Flow Matching approach of Lipman et al. (2023) circumvents this by using a regression objective to train a CNF, which avoids simulations during training,
@@
where $p_{t}(\cdot|Y)$ is a time dependent conditional density path associated to the conditional vector field $v_{t}(\cdot|Y)$ and both $p_{t}(\cdot|Y)$ and $v_{t}(\cdot|Y)$ are known by construction. This regression objective is in certain cases similar to the score matching objective (Hyvärinen, 2005; Vincent, 2011), which is used to train diffusion models based on stochastic differential equations, introduced by Song et al. (2021).
@@
Flow Matching algorithms have been successfully used in many different applications that benefit from efficient sampling, such as text-to-speech (Guo et al., 2024) and text-to-image (Yang et al., 2025; Esser et al., 2024) settings, the production of novel molecular or protein structures (Dunn and Koes, 2024; Bose et al., 2024) or the construction of surrogate models in high energy physics (Bieringer et al., 2024). It has also been adjusted theoretically to different settings. Atanackovic et al. (2024) adapted Flow Matching to the case of interacting particles, Gat et al. (2024) explored the discrete setting. Chen and Lipman (2024) generalized the Euclidean setting to the Riemannian setting, allowing for more general geometries. Kerrigan et al. (2024) extended Flow Matching to function spaces. However, the statistical properties have only recently been studied by Fukumizu et al. (2025) in the Wasserstein $2$ distance. They do not use the exact setting of Lipman et al. (2023), but rather introduce stopping times that depend on the number of samples, enabling the transfer of methods known from the statistical analysis of diffusion models.
@@
In this work, we demonstrate that Flow Matching in the setting of Lipman et al. (2023) is closely related to the classical kernel density estimation (KDE). This connection allows us to analyze Flow Matching from a new perspective. First, we show that the motivation of Flow Matching also holds for its empirical counterparts where we replace the unknown target density $p^{*}$ in (1) with the empirical measure of the observations. For sufficiently large network classes, the resulting generative algorithm coincides exactly with a kernel density estimator, where the kernel is given by the density of the latent distribution. We show convergence rates for the kernel density estimator in Wasserstein $1$ distance in a setting that allows for the Gaussian kernel which is the most popular choice in Flow Matching applications. Our rates coincide with the minimax optimal rates from Niles-Weed and Berthet (2022) up to logarithmic factors. We then transfer our analysis to Flow Matching using neural network classes of parameterized vector fields. This particularly shows that even in case of perfect approximation, which corresponds to heavy overparameterization in practice, the Flow Matching estimator does not deteriorate to the empirical measure, but to an estimator that is minimax optimal. Put another way, Flow Matching can be, up to constants, at least as good as a KDE even without the presence of network-specific structures.
@@
Unlike Fukumizu et al. (2025), we do not build on the similarities to diffusion models. Separating the error of the kernel density estimator allows us to analyse empirical risk minimization without the need for Chernoff-type bounds. This avoids the problems pointed out by Yakovlev and Puchkin (2025) in the analysis of diffusion models. Overall, the analysis of Flow Matching is more delicate, since Girsanov’s theorem does not apply and thus the strategy of Chen et al. (2023c) cannot be used. Therefore, our bound depends exponentially on the Lipschitz constant of the vector field. This is one of the reasons for the difficulties in the proof of Fukumizu et al. (2025) [^2] The bounds of the integral in the exponent of the first equation on p. 14 (in the ArXiV version) are not correct. The separation of the outer integral does not imply the same separation of the integral in the exponent. The correct factor in the integral is $\exp(2\int_{T_{0}}^{t_{j}}L_{u}du)$, where $L_{u}$ is the Lipschitz constant of the approximated vector field, which depends on $T_{0}$ and is thus not bounded by a universal constant.. The main theoretical novelty of this work compared to standard approaches in the statistical literature on diffusions and Flow Matching is to use the KDE as a reference model for the empirical risk minimizer. Therefore, we do not recover the classical trade-off in network size. Since in general, assuming that the Lipschitz constant of the vector field is bounded with a fixed constant requires an assumption on the unknown distribution (Kunkel, 2025), our approach requires significantly less assumptions on the unknown distribution. Our analysis allows us to show convergence rates in Wasserstein $1$ distance in the case where overparameterized neural networks are used for the vector field.
While these rates are minimax optimal up to a logarithmic constant, they suffer from the curse of dimensionality. In case $\mathbb{P}^{*}$ is supported on a lower-dimensional manifold, we show that, on a small tube around the manifold, the mass of the generated distribution concentrates according to the unknown distribution much faster, depending on the intrinsic dimension and thus circumventing the curse of dimensionality. This result again applies to both, the KDE and the Flow Matching estimator. For the former, our results are the first that do not require a structural adaption of the KDE to the manifold itself. Therefore, they contribute to the statistical understanding of the KDE under the manifold hypothesis previously studied by Divol (2022); Berenfeld and Hoffmann (2021).
@@
Lipman et al. (2023) use a fixed latent distribution. Similar approaches for flows between two possibly unknown distributions $\mathbb{P}$ and $\mathbb{Q}$ are studied by Tong et al. (2024), Liu et al. (2022) and Albergo and Vanden-Eijnden (2023). Tong et al. (2024) also generalize the mentioned methods. Gao et al. (2024) prove a suboptimal rate of convergence for certain forms of Flow Matching not exactly conforming to the construction on Lipman et al. (2023) in the Wasserstein $2$ distance. Benton et al. (2024) analyzed Flow Matching excluding the approximation error by imposing assumptions on the covariance that lead to global Lipschitz bounds on the vector field. They also focus on different constructions than Lipman et al. (2023). Gong et al. (2025) study the properties of ReLU networks to approximate a vector field corresponding to higher order trajectories.
As Lipman et al. (2023) point out, Flow Matching is closely related to diffusion models. For an overview of generative diffusion models, see Cao et al. (2024). Indeed, even in cases that are not constructed to be consistent with diffusion models, the approximation of a score function has similar properties to the approximation of the Flow Matching vector field. Fukumizu et al. (2025) build up on this connection. The statistical properties of score matching are an area of ongoing research, see for example Chen et al. (2023a, b, c); Oko et al. (2023); Tang and Yang (2024); Azangulov et al. (2024); Zhang et al. (2024); Yakovlev and Puchkin (2025). Marzouk et al. (2023) study the statistical properties of CNFs trained by likelihood maximization.
The kernel density or Parzen-Rosenblatt estimator (Rosenblatt, 1956; Parzen, 1962) is the classical method for estimating a smooth density. For an overview in the univariate case, see Tsybakov (2009), Chapter 1.2 or Devroye and Lugosi (2001) and in the multivariate case see Scott (1992). Typically, the kernel density estimator is analyzed using the mean squared error or the $L_{1}$ distance. Despite its longstanding use, kernel density estimation is still the subject of ongoing research, for example in density estimation on unknown manifolds (Berry and Sauer, 2017; Berenfeld and Hoffmann, 2021; Divol, 2022; Wu and Wu, 2022). In particular, Divol (2022) uses the Wasserstein metric to evaluate the performance of a kernel density estimator, however he needs to assume properties of the kernel that are not satisfied by standard choices such as the Gaussian kernel.
@@
In Section˜2 we introduce the problem setting mathematically and recall all the definitions needed to define the conditional Flow Matching objective from Lipman et al. (2023). We introduce the empirical counterparts, validate the empirical conditional Flow Matching objective, and portray the connection to the kernel density estimator. In Section˜3 we first separate the error corresponding to the kernel density estimator and the vector field approximation. We then proceed to bound the former error in Section˜3.1. Incorporating theory of network approximation, we bound the latter error and obtain a rate of convergence for the entire Flow Matching procedure in Section˜3.2. Subsequently, we show in Section˜3.3 that the rate can be improved if the unknown distribution is supported on a lower-dimensional manifold. Finally, we illustrate our results in Section˜4. All proofs are deferred to Section˜5.
@@ 1.1 Some notations
We write $X\sim p$ as a synonym for $X\sim\mathbb{P}$ in case $\mathbb{P}$ has a density $p$ with respect to the Lebesgue measure and $\mathbb{P}$ is the distribution of the random variable $X$. By $\lesssim$ we denote $a\leq c\cdot b,$ where $a,b\in\mathbb{R}$ and $c$ is a constant independent of $n$. For $d\in\mathbb{N}$, we use the Euclidean norm $|\cdot|$ on $\mathbb{R}^{d}$ and use $\|\cdot\|$ for associated operator norms. For a set $A\subset\mathbb{R}^{d}$, we denote the tubular neighborhood of the set $A$ with radius $r>0$ with $B(r,A)$. The supremum norm and the $L_{1}$-norm of a function $f\colon\Omega\rightarrow\mathbb{R}^{d^{\prime\prime}}$ with $d^{\prime},d^{\prime\prime}\in\mathbb{R}$ and $\Omega\subset\mathbb{R}^{d^{\prime}}$ is defined as
@@
For $d^{\prime\prime}=1,\alpha\in(0,1]$ and $\Lambda>0$, the Besov ball $B_{1,\infty}^{\alpha}(\Lambda,\infty)$ is defined as
$B_{1,\infty}^{\alpha}(\Lambda,\Omega)\coloneqq\big\{f\in L_{1}(\Omega)\colon|f|_{B_{1,\infty}^{\alpha}(\Omega)}\coloneqq\sup_{t>0}t^{-\alpha}\omega_{1}(f,t)_{1}\leq\Lambda\big\}$ (2)
@@
The difference is set to zero if $x+h\notin\Omega.$ Replacing $|f|_{B_{1,\infty}^{\alpha}(\Omega)}<\Lambda$ with $|f|_{B_{1,\infty}^{\alpha}(\Omega)}<\infty$ in (2), we obtain the Besov space $B_{1,\infty}^{\alpha}(\Omega)$. On $B_{1,\infty}^{\alpha}(\Omega)$ we define the norm $\|f\|_{B_{1,\infty}^{\alpha}(\Omega)}=\|f\|_{L_{1}(\Omega)}+|f|_{B_{1,\infty}^{\alpha}(\Omega)}$. Compared to more commonly used smoothness classes, such as Hölder classes, the Besov class defined above only imposes integrated and not pointwise smoothness. This allows for local irregularities like jumps.
@@ 2 Empirical Flow Matching
The ODE (3) corresponds to the Lagrangian description (in terms of particle trajectories) of the conservation of mass formula (Villani, 2008, p.14). The change of variables formula links it to the Eulerian description: A necessary and sufficient condition for $v_{t}$ to generate $p_{t}$ is
@@
by Lipman et al. (2023).
They have consider a probability path of the form
@@
$v_{t}(x)=\int v_{t}(x|y)\frac{p_{t}(x|y)p^{*}(y)}{p_{t}(x)}dy.$ (7)
In this setting Lipman et al. (2023) show that the Flow Matching objective (5) is equivalent to the conditional Flow Matching objective
@@
With this modification we recover the sufficient condition for $v^{n}_{t}$ to generate $p^{n}_{t}$ analogously to (Lipman et al., 2023, Theorem 1).
@@
The motivation for the conditional Flow Matching objective (8) is the equivalence to the unconditioned Flow Matching objective with respect to the optimizing arguments (Lipman et al., 2023, Theorem 2). Using the empirical counterparts, this still holds true.
@@
Using $p^{n}_{t}$ and $v^{n}_{t}$ from (9), and a class of parameterized vector fields $\mathcal{N},$ which is constructed such that the minimal arguments exists, we have that
$\underset{\hat{v}\in\mathcal{N}}{\mathrm{argmin}}\int_{0}^{1}\mathbb{E}_{X_{t}\sim p^{n}_{t}}\big[|\hat{v}_{t}(X_{t})-v^{n}_{t}(X_{t})|^{2}\big]\;\mathrm{d}t=\underset{\hat{v}\in\mathcal{N}}{\mathrm{argmin}}\int_{0}^{1}\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\tilde{X}_{t}\sim p_{t}(\cdot|X_{i}^{*})}\big[|\hat{v}_{t}(\tilde{X}_{t})-v_{t}(\tilde{X}_{t}|X_{i}^{*})|^{2}\big]\;\mathrm{d}t.$
@@
is justified theoretically. Compared to Gao et al. (2024), we use $v_{t}(\cdot|X_{i}^{*})$ directly as proposed by Lipman et al. (2023) and stick to the entire time interval. Note that for simplicity we do not sample $t.$
From minimizing (10) in $\tilde{v}\in\mathcal{N},$ we obtain an optimal argument $\hat{v}$. Solving the ODE (3) using $\hat{v}$, we obtain a flow $\hat{\psi}_{t}$, i.e. $\hat{\psi}$ is given by
[…]
@@
we recover the same result like in (Lipman et al., 2023, Theorem 3):
@@
Set $\sigma_{\min}>0.$ To flow from $p^{n}_{0}(x)=\frac{1}{n}\sum_{i=1}^{n}K(x)=K(x)$ to $p^{n}_{1}(x)=\frac{1}{n\sigma^{d}_{\min}}\sum_{i=1}^{n}K(\frac{x-X_{i}^{*}}{\sigma_{\min}}),$ we can choose any differentiable functions $\sigma_{t}$ and $\mu_{t}$ such that
@@
$p^{n}_{1}(x)=\frac{1}{n\sigma^{d}_{\min}}\sum_{i=1}^{n}K\Big(\frac{x-X_{i}^{*}}{\sigma_{\min}}\Big),$ (14)
where the kernel is given by the latent distribution $\mathbb{U}$. Choosing $\mathbb{U}=\mathcal{N}_{d}(0,1)$, i.e., we consider the $d$-dimensional Gaussian kernel $K(x)=(2\pi)^{-d/2}\exp(-|x|^{2}/2)$, yields the proposed flow from Lipman et al. (2023, Section 4). Moreover, considering general kernels is in line with methods by Tong et al. (2024), Liu et al. (2022) and Albergo and Vanden-Eijnden (2023), which transform an unknown distribution to another.